OpenStudy (anonymous):

The product of 9 and a number is 12 less than 3 times that number. What is the number? 9*n-12=3*n

OpenStudy (anonymous):

My guess is 2... I just need someone to walk me through it again, and/or verify my answer.

OpenStudy (er.mohd.amir):

9*n-12=3*n 9*n-3*n=12 6*n=12 n=12/6 n=2